# Driving with Prior Maps: Unified Vector Prior Encoding for Autonomous Vehicle Mapping

### v2

@@ Abstract
High-Definition Maps (HD maps) are essential for the precise navigation and decision-making of autonomous vehicles, yet their creation and upkeep present significant cost and timeliness challenges. The online construction of HD maps using on-board sensors has emerged as a promising solution; however, these methods can be impeded by incomplete data due to occlusions and inclement weather. This paper proposes the PriorDrive framework to addresses these limitations by harnessing the power of prior maps, significantly enhancing the robustness and accuracy of online HD map construction. Our approach integrates a variety of prior maps, such as OpenStreetMap’s Standard Definition Maps (SD maps), outdated HD maps from vendors, and locally constructed maps from historical vehicle data. To effectively encode this prior information into online mapping models, we introduce a Hybrid Prior Representation (HPQuery) that standardizes the representation of diverse map elements. At the core of PriorDrive is the Unified Vector Encoder (UVE), which employs a dual encoding mechanism to process vector data. The intra-vector encoder captures fine-grained local features, while the inter-vector encoder integrates global context. Furthermore, we propose a segment-level and point-level pre-training strategy that enables the UVE to learn the prior distribution of vector data, thereby improving the encoder’s generalizability and performance. Through extensive testing on the nuScenes dataset, we demonstrate that PriorDrive is highly compatible with various online mapping models and substantially improves map prediction capabilities. The integration of prior maps through the PriorDrive framework offers a robust solution to the challenges of single-perception data, paving the way for more reliable autonomous vehicle navigation.
Introduction
@@
High-Definition Maps (HD maps) are essential for the accurate navigation and decision-making of autonomous vehicles, providing detailed vectorized representations of road elements such as lane dividers, pedestrian crossings, and road boundaries (Elghazaly et al. 2023; Gao et al. 2020). Despite their significance, the traditional methods for creating and maintaining HD maps are often costly and labor-intensive. These methods can result in outdated maps that struggle to keep pace with the rapidly changing urban environments. In response, there is increasing interest in online HD map construction (Qiao et al. 2023; Ding et al. 2023; Zhang et al. 2024a), where maps are generated in real-time using on-board sensors, such as panoramic RGB cameras and LiDAR point clouds. Although this approach reduces costs, it also introduces new challenges, particularly concerning incomplete and error-prone data caused by environmental occlusions or inclement weather conditions.
@@
In this paper, we introduce PriorDrive, a novel framework designed to overcome these challenges by leveraging prior maps. Specifically, PriorDrive incorporates three types of prior maps: Standard Definition Maps (SD maps), existing but outdated HD maps, and online historical prediction maps, as illustrated in Figure LABEL:fig:priors. SD maps provide essential long-distance centerline skeletons, which are crucial for maintaining road-level connectivity. Existing HD maps (HD map-EX), despite their high accuracy, often fail to reflect current road conditions due to infrequent updates. Nevertheless, these maps offer valuable prior information for HD map construction. Additionally, historical prediction maps, which are often overlooked, can provide insights into current road conditions by reflecting the latest observations from previous perceptions. As vehicles traverse the same locations multiple times under varying conditions, the accumulated information can be used to enhance the quality of the current local map through iterative refinement.
@@
Recent studies (Xiong et al. 2023; Jiang et al. 2024; Sun et al. 2023) have attempted to use prior maps to supplement perception data, but they face several limitations. For instance, NMP (Xiong et al. 2023) relies on a neural prior that cannot take advantage of existing offline maps and can only be updated through model training, leading to poor transferability to new environments. P-MapNet (Jiang et al. 2024) incorporates vector maps into the pipeline by converting them into rasterized representations, which are lossy and redundant, lacking the detail necessary to capture vectorized instance-level information and express the type and direction of map elements effectively. MapEX (Sun et al. 2023), on the other hand, focuses on existing HD maps but simulates outdated maps by introducing artificial offsets and erasing map elements, an approach that risks leaking ground truth data and fails to accurately represent the differences between simulated and actual outdated HD maps. These actual outdated maps, due to their low update frequency, struggle to reflect real-time changes in road structures, making them unreliable for current navigation needs.
@@
Given that various maps are typically stored as vectors due to their low redundancy and universal applicability, it’s crucial to develop a robust vector encoding model. However, designing such a model is challenging due to the variable length, sparsity, and discreteness of vector data. To address this and enhance current perception results by integrating diverse vector prior maps into various mapping models, we propose a unified vector encoder (UVE). The UVE can encode a wide range of vector data types, and we introduce pre-training paradigms that allow the model to fully learn the prior distribution of vector maps. Our experiments on the nuScenes dataset demonstrate that our PriorDrive can be seamlessly applied to various online mapping models, resulting in significant performance improvements, including a +2.2 mIoU for HDMapNet and a +4.2 mAP for MapTRv2.
@@
  • We introduce a unified vector encoder (UVE) that effectively encodes various vector data through a dual encoding mechanism: the Intra-Vector Encoder captures fine local features, while the Inter-Vector Encoder fuses global context information.
  • We propose a pre-training paradigm for vector data that learns the prior distribution of vector data by adding Gaussian noise or mask at the segment and point levels and reconstructing the entire vector map.
@@
  • We propose a hybrid prior representation (HPQuery) to represent all elements and a PriorDrive framework with vector prior maps to addresses the limitations of single-perception. Our comprehensive evaluation on nuScenes dataset demonstrates that our PriorDrive significantly enhances the performance of online mapping models.
Related Work
Online Vectorized HD Map Construction
Recent advances in HD map construction have shifted from traditional offline methods based on SLAM techniques (Shan and Englot 2018; Shan et al. 2020) to more dynamic approaches that utilize vehicle-mounted sensors for online HD map generation (Xie, Pang, and Wang 2023; Chen, Deng, and Furukawa 2024; Liu et al. 2024a). This shift not only reduces costs but also ensures that maps reflect up-to-date road conditions. Some approaches (Li et al. 2021, 2022) treat HD map construction as a segmentation task, predicting pixel-level rasterized maps that require complex post-processing for vectorization. To address this limitation, recent work (Hu et al. 2024; Zhang et al. 2024b; Chen et al. 2024) has focused on developing end-to-end frameworks for vectorized map learning.
@@
VectorMapNet (Liu et al. 2022) introduced a coarse-to-fine, two-stage network leveraging keypoint representations. MapTR (Liao et al. 2022) advanced this concept by modeling point sets in a permutation-equivalent manner using a DETR-like (Carion et al. 2020) one-stage network. HIMap (Zhou et al. 2024) proposed a hybrid representation learning pipeline to capture both point-level and element-level information. MapQR (Liu et al. 2024c) introduced a scatter-and-gather query mechanism for precise detection of content and location information, while MGMap (Liu et al. 2024b) used mask-guided features to refine lane shape representations with enhanced feature detail. However, these methods rely solely on single-source perception data from vehicle-mounted sensors, limiting their effectiveness in challenging scenarios such as occlusion or inclement weather conditions.
Online Mapping Based on Prior Maps
@@
Recent research has explored the integration of prior maps to enhance the performance of online mapping models. P-MapNet (Jiang et al. 2024) encodes Standard Definition Maps (SD map) as an additional conditional branch and employs a masked autoencoder to capture the prior distribution of HD maps. NMP (Xiong et al. 2023) introduces a global neural map prior that self-updates, thereby improving the performance of local map inference. MapEX (Sun et al. 2023) categorizes existing maps into three types and refines the query-based map estimation model’s matching algorithm. Despite these advances, many studies overlook the potential of online historical prediction maps as a source of prior information. Our work addresses this gap by leveraging various existing prior maps in their native vector form to enhance current perception data.
Pretrained Methods Based on Mask Modeling
@@
In the fields of Natural Language Processing and Computer Vision, masked pre-training has proven to be an effective strategy for self-supervised representation learning. In NLP, models like BERT (Devlin et al. 2019; Liu et al. 2019; Lan et al. 2019) use masked language modeling to predict randomly masked tokens within a bidirectional text context. Similarly, methods such as MAE (He et al. 2021; Li et al. 2023; Liu et al. 2023) mask random patches of input images and reconstruct them based on the remaining unmasked patches. Diverging from these approaches, which are predominantly tailored for text and image data, we introduce a pre-training paradigm specifically designed for vector data. Our method aims to learn the prior distribution of vector representations, offering a new avenue for pre-training in the context of vectorized map data.
Proposed Method: PriorDrive
@@
Figure 1: Our Integration Method. For query-based models, we enable interaction between the prior features and queries at both the instance-level and point-level, ensuring a more nuanced integration. Hierarchical query comes from MapTRv2 (Liao et al. 2023). For the other models, we integrate the prior features directly with the BEV features.
Problem Formulation
@@
The goal of online mapping is to construct a local HD map using on-board sensor observations, such as the surrounding camera images. The online mapping process typically involves two main modules: the Bird’s-Eye View (BEV) encoder and the map decoder. The BEV encoder generally includes a Perspective View (PV) feature extractor and a module for transforming PV features into BEV features. The map decoder usually comprises a Fully Convolutional Network (FCN) (Shelhamer, Long, and Darrell 2015) or a query-based module similar to DETR (Carion et al. 2020), which outputs either rasterized perception results or vectorized map elements.
@@
We proposed four methods in total to integrate prior map features into online mapping models. For the methods without a learnable query $Q$, such as HDMapNet (Li et al. 2021), we directly reshape the prior feature $f_{prior}$ and use $Deconv$ upsampling to match the shape of the BEV feature $f_{BEV}$. Then connect it with the BEV feature $f_{BEV}$ and aligned by $Conv$, as shown in Figure 1 left.
@@
here, $f_{bev}^{*}$ represents the BEV feature after incorporating the prior map information, which can be input into the map decoder $D_{map}$ to obtain the predicted results of map elements.
For methods based on learnable query $Q$, such as the vectorized model MapTR series (Liao et al. 2023), we propose HPQuery of three operations: addition, replacement, and concatenation, as shown is Figure 1 right. Specifically, $Q$ is typically composed of two parts, $Q=\{q_{ij}\}_{i=0,j=0}^{m,n}=\{q_{i}^{ins}+q_{j}^{pt}\}_{i=0,j=0}^{m,n}$, $q_{i}^{ins}$ and $q_{i}^{pt}$ representing the learnable queries at the instance-level and point-level respectively. The features $f_{prior}$ of prior maps are also composed of instance-level features and point-level features, $f_{prior}=\{(f_{i}^{ins},f_{j}^{pt})\}_{i=0,j=0}^{m^{\prime},n^{\prime}}$. So we interact prior features and queries at the instance level and click, respectively. The HPQuery of addition, replacement, and concatenation operations can be expressed as:
@@
Architecture of UVE
Inspired by BERT (Devlin et al. 2019) on text-related tasks, we propose a unified vector encoder (UVE) by analogizing vector points to words and vector map elements to sentences (see Figure 2). And we provide a pre-training strategy to improve the encoding and noise reduction ability of UVE.
@@
As shown in Figure 2, for the given prior map $M_{prior}$, we first construct the input for UVE, termed as Hybrid Prior Embedding (HPE). We use the method from (Tancik et al. 2020) for obtaining point position embedding and direction embedding for both $(x,y)$ and $(v_{x},v_{y})$. We concatenate these two embeddings as the point-level embedding. To obtain the feature representation of the entire vector, we add a special [VEC] token at the beginning of each vector and use the embedding of this token as the instance-level embedding. Additionally, we introduce learnable instance embedding and learnable type embedding to distinguish vector instances. To ensure the order of points, we also introduce 2D learnable position embedding. Finally, these embeddings are aggregated to form the HPE, which then serves as the input embedding for UVE.
Intra and Inter-Vector Encoder.
Due to the limited information content of vector points themselves, learning the concept of vector instances is very challenging for the model, requiring more interactions among the points within the vectors to enhance each point’s perceptual abilities towards others in the same vector. As a result, the UVE encoder utilizes $M$-layer intra-vector attention encoders and $N$-layer inter-vector attention encoders, using an attention masking mechanism to facilitate feature interactions within and between different vector instances. Through this dual encoding mechanism, UVE can deeply understand and represent vector data, providing strong support for key tasks such as online HD maps construction.
@@
Pre-training UVE: Position Modeling
@@
where $P=\{(x_{i},y_{i})\}_{i=0}^{k}$ represent all the point in the $M_{org}$.
Experiments
Experimental Settings

### v3

@@ Abstract
High-Definition Maps (HD maps) are essential for the precise navigation and decision-making of autonomous vehicles, yet their creation and upkeep present significant cost and timeliness challenges. The online construction of HD maps using on-board sensors has emerged as a promising solution; however, these methods can be impeded by incomplete data due to occlusions and inclement weather. This paper proposes the PriorDrive framework to addresses these limitations by harnessing the power of prior maps, significantly enhancing the robustness and accuracy of online HD map construction. Our approach integrates a variety of prior maps, such as OpenStreetMap’s Standard Definition Maps (SD maps), outdated HD maps from vendors, and locally constructed maps from historical vehicle data. To effectively encode this prior information into online mapping models, we introduce a Hybrid Prior Representation (HPQuery) that standardizes the representation of diverse map elements. At the core of PriorDrive is the Unified Vector Encoder (UVE), which employs hybrid prior embedding and a dual encoding mechanism to process vector data. Furthermore, we propose a segment-level and point-level pre-training strategy that enables the UVE to learn the prior distribution of vector data, thereby improving the encoder’s generalizability and performance. Through extensive testing on the nuScenes, Argoverse 2 and OpenLane-V2, we demonstrate that PriorDrive is highly compatible with various online mapping models and substantially improves map prediction capabilities. The integration of prior maps through the PriorDrive framework offers a robust solution to the challenges of single-perception data, paving the way for more reliable autonomous vehicle navigation.
Figure 1: Overview of PriorDrive. Our unified vector encoder (UVE) directly encodes diverse vector prior maps and seamlessly integrates them into existing online mapping frameworks. This integration enhances the final predictions, making them more complete and accurate than those generated without prior information.
footnotetext: Work done during the internship at Amap, Alibaba Group.footnotetext: Equal contribution.footnotetext: Corresponding author.
1 Introduction
@@
High-Definition Maps (HD maps) are essential for the accurate navigation and decision-making of autonomous vehicles, providing detailed vectorized representations of road elements such as lane dividers, pedestrian crossings, and road boundaries [6, 8]. Despite their significance, the traditional methods for creating and maintaining HD maps are often costly and labor-intensive. These methods can result in outdated maps that struggle to keep pace with the rapidly changing urban environments. In response, there is increasing interest in online HD map construction [27, 5, 38], where maps are generated in real-time using on-board sensors, such as panoramic RGB cameras and LiDAR point clouds. Although this approach reduces costs, it also introduces new challenges, particularly concerning incomplete and error-prone data caused by environmental occlusions or inclement weather conditions.
@@
In this paper, we introduce PriorDrive, a novel framework designed to overcome these challenges by leveraging prior maps. Specifically, PriorDrive incorporates three types of prior maps: Standard Definition Maps (SD maps), existing but outdated HD maps, and online historical prediction maps, as illustrated in Figure 1. SD maps provide essential long-distance centerline skeletons, which are crucial for maintaining road-level connectivity. Existing HD maps (HD map-EX), despite their high accuracy, often fail to reflect current road conditions due to infrequent updates. Nevertheless, these maps offer valuable prior information for HD map construction. Additionally, historical prediction maps, which are often overlooked, can provide insights into current road conditions by reflecting the latest observations from previous perceptions. As vehicles traverse the same locations multiple times under varying conditions, the accumulated information can be used to enhance the quality of the current local map through iterative refinement.
@@
Recent studies [35, 13, 29] have attempted to use prior maps to supplement perception data, but they face several limitations. For instance, NMP [35] relies on a neural prior that cannot take advantage of existing offline maps and can only be updated through model training, leading to poor transferability to new environments. P-MapNet [13] incorporates vector maps into the pipeline by converting them into rasterized representations, which are lossy and redundant, lacking the detail necessary to capture vectorized instance-level information and express the type and direction of map elements effectively. MapEX [29], on the other hand, focuses on existing HD maps but simulates outdated maps by introducing artificial offsets and erasing map elements, an approach that risks leaking ground truth data and fails to accurately represent the differences between simulated and actual outdated HD maps. These actual outdated maps, due to their low update frequency, struggle to reflect real-time changes in road structures, making them unreliable for current navigation needs.
In addition to constructing accurate online HD maps, the autonomous driving scenario requires understanding the topology in road scenes (i.e., lane connectivity and correspondence with traffic elements), which is called topology reasoning task. Since connected SD maps include crucial information for road topology, we also apply PriorDrive to this task to enhance the performance of topology reasoning.
@@
Given that various maps are typically stored as vectors due to their low redundancy and universal applicability, it’s crucial to develop a robust vector encoding model. However, designing such a model is challenging due to the variable length, sparsity, and discreteness of vector data. To address this and enhance current perception results by integrating diverse vector prior maps into various mapping models, we propose a unified vector encoder (UVE). The UVE can encode a wide range of vector data types, and we introduce pre-training paradigms that allow model to fully learn the prior distribution of vector maps. Our experiments on nuScenes, Argoverse 2 and OpenLane-V2 demonstrate that our PriorDrive is plug-and-play and can be seamlessly applied to various online mapping models and vector tasks, resulting in significant performance improvements.
@@
  • We introduce a unified vector encoder that effectively captures diverse vector attributes and enables encoding of variable-length vectors into fixed-length features through Hybrid prior embedding and dual encoding mechanism.
  • We first proposed a pre-training paradigm for vector data, which can effectively learn the prior distribution of vector data and help reduce noise in historical maps through point-level and segment-level pre-training.
@@
  • We propose a hybrid prior representation (HPQuery) to represent all elements and a PriorDrive framework with vector prior maps to addresses the limitations of single-perception. Our comprehensive evaluation demonstrates that PriorDrive is plug-and-play and significantly enhances the performance of online mapping models.
2 Related work
2.1 Online vectorized HD map construction
@@
Online vectorized HD map construction was initially treated as a segmentation task [15, 18]. To construct vectorized HD map, HDMapNet [15] requires complex post-processing of pixel-level rasterized maps. VectorMapNet [25] introduced a coarse-to-fine, two-stage network leveraging keypoint representations. MapTR [19] advanced this concept by modeling point sets in a permutation-equivalent manner using a DETR-like [2] one-stage network. Subsequent studies [41, 39, 36, 9, 12, 40, 3] have introduced various insightful approaches for further improvement. However, these methods rely solely on single-source perception data from vehicle-mounted sensors, limiting their effectiveness in challenging scenarios such as occlusion or inclement weather conditions.
The release of OpenLane-V2 [31] dataset has sparked growing interest in topology reasoning within driving scenes. This task focuses on recognizing the topologies among lanes and lanes with traffic elements. TopoNet [17] leverages GNN to update lane representation and lane topology. TopoMLP [34] utilizes PETR [24] for centerline detection and employs simple MLPs to predict the relationships. In this paper, we also apply PriorDrive to topological inference tasks, introducing interconnected SD Map as an additional input to support topology reasoning.
2.2 Online mapping based on prior maps
@@
Recent research has explored the integration of prior maps to enhance the performance of online mapping models. P-MapNet [13] encodes Standard Definition Maps (SD map) as an additional conditional branch and employs a masked autoencoder to capture the prior distribution of HD maps. NMP [35] introduces a global neural map prior that self-updates, thereby improving the performance of local map inference. MapEX [29] categorizes existing maps into three types and refines the query-based map estimation model’s matching algorithm. Despite these advances, many studies overlook the potential of online historical prediction maps as a source of prior information. Our work addresses this gap by leveraging various existing prior maps in their native vector form to enhance current perception data.
2.3 Pretrained methods based on mask modeling
@@
In the fields of NLP and CV, masked pre-training has proven to be an effective strategy for self-supervised representation learning. In NLP, models like BERT [4, 23, 14] use masked language modeling to predict randomly masked tokens within a bidirectional text context. Similarly, methods such as MAE [11, 16, 21] mask random patches of input images and reconstruct them based on the remaining unmasked patches. Diverging from these approaches, which are predominantly tailored for text and image data, we introduce a pre-training paradigm specifically designed for vector data. Our method aims to learn the prior distribution of vector representations, offering a new avenue for pre-training in the context of vectorized data.
3 Proposed method: PriorDrive
@@
Figure 2: Our Integration Method. For query-based models, we enable interaction between the prior features and queries at both instance-level and point-level, ensuring a more nuanced integration. Hierarchical query comes from MapTRv2 [20]. For the other models, we integrate prior features directly with the BEV features.
3.1 Problem formulation
@@
The goal of online mapping is to construct a local HD map using on-board sensor observations, such as the surrounding camera images. The online mapping process typically involves two main modules: the Bird’s-Eye View (BEV) encoder and the map decoder. The BEV encoder generally includes a Perspective View (PV) feature extractor and a module for transforming PV features into BEV features. The map decoder usually comprises a Fully Convolutional Network [28] or a query-based module similar to DETR [2], which outputs either rasterized perception results or vectorized map elements.
@@
We proposed four methods in total to integrate prior map features into online mapping models. For the methods without a learnable query $Q$, such as HDMapNet [15], we directly reshape the prior feature $f_{prior}$ and use $Deconv$ upsampling to match the shape of the BEV feature $f_{bev}$. Then connect it with the BEV feature $f_{bev}$ and aligned by $Conv$, as shown in Figure 2 left.
@@
here, $f_{bev}^{*}$ represents BEV feature after incorporating the prior map information, which can be input into the map decoder $D_{map}$ to obtain the predicted results of map elements.
For methods based on learnable query $Q$, such as the vectorized model MapTR series [20], we propose HPQuery of three operations: addition, replacement, and concatenation, as shown is Figure 2 right. Specifically, $Q$ is typically composed of two parts, $Q=\{q_{ij}\}_{i=0,j=0}^{m,n}=\{q_{i}^{ins}+q_{j}^{pt}\}_{i=0,j=0}^{m,n}$, $q_{i}^{ins}$ and $q_{i}^{pt}$ representing the learnable queries at the instance-level and point-level respectively. The features $f_{prior}$ of prior maps are also composed of instance-level features and point-level features, $f_{prior}=\{(f_{i}^{ins},f_{j}^{pt})\}_{i=0,j=0}^{m^{\prime},n^{\prime}}$. So we interact prior features and queries at the instance level and click, respectively. The HPQuery of addition, replacement, and concatenation operations can be expressed as:
@@
3.2 Architecture of UVE
Inspired by BERT [4] on text-related tasks, we propose a unified vector encoder (UVE) by analogizing vector points to words and vector map elements to sentences (see Figure 3). And we provide a pre-training strategy to improve the encoding and noise reduction ability of UVE.
@@ 3.2 Architecture of UVE
As shown in Figure 3, for the given prior map $M_{prior}$, we first construct the input for UVE, termed as Hybrid Prior Embedding (HPE). We use the method from [30] for obtaining point position embedding and direction embedding for both $(x,y)$ and $(v_{x},v_{y})$. We concatenate these two embeddings as the point-level embedding. To obtain the feature representation of the entire vector, we add a special [VEC] token at the beginning of each vector and use the embedding of this token as the instance-level embedding. Additionally, we introduce learnable instance embedding and learnable type embedding to distinguish vector instances. To ensure the order of points, we also introduce 2D learnable position embedding. Finally, these embeddings are aggregated to form the HPE, which then serves as the input embedding for UVE.
Dual encoding mechanism.
Due to the limited information content of vector points themselves, learning the concept of vector instances is very challenging for the model, requiring more interactions among the points within the vectors to enhance each point’s perceptual abilities towards others in the same vector. As a result, the UVE encoder utilizes $M$-layer intra-vector attention encoders and $N$-layer inter-vector attention encoders, using an attention masking mechanism to facilitate feature interactions within and between different vector instances.
@@
3.3 Pre-training UVE: Position modeling
@@
where $P=\{(x_{i},y_{i})\}_{i=0}^{k}$ represent all the point in $M_{org}$.
4 Experiments
4.1 Experimental settings
HDMapNet [15] C Effi-B0 30 IoU 18.7 40.6 39.5 32.9 -
# SD map Prior (Ours) C Effi-B0 30 IoU 22.2 42.0 41.5 35.2 -
# Online Local Map Prior (Ours) C Effi-B0 30 IoU 23.1 41.4 42.0 35.5 -
# HD map-EX* Prior (Ours) C Effi-B0 30 IoU 21.2 42.3 44.9 36.1 -
MapTR [19] C R50 24 AP 41.2 49.5 51.1 47.3 -
# P-MapNet [13] C R50 24 AP 43.7 50.9 53.5 49.4 -
StreamMapNet [37] C R50 24 AP 60.4 61.9 58.9 60.4 -
# SQD-MapNet [32] C R50 24 AP 63.0 65.5 63.3 63.9 -
MGMap [22] C R50 24 AP 61.8 65.0 67.5 64.8 -
MapTRv2 [20] C R50 24 AP 59.8 62.4 62.4 61.5 38.6
# SD map Prior (Ours) C R50 24 AP 61.7 65.5 66.4 64.5 42.7
# Online Local Map Prior (Ours) C R50 24 AP 63.6 67.3 66.2 65.7 44.7
# HD map-EX* Prior (Ours) C R50 24 AP 61.5 65.6 68.3 65.1 44.3
Table 1: Performance of various online mapping models with different prior maps on nuScenes within a $60m\times 30m$ perception range. “C” denotes camera input. The SD map is sourced from OpenStreetMap [10]. The online local map refers to the historical prediction results. As the nuScenes dataset lacks existing HD maps, we followed the approach of MapEX [29] to create an HD map-EX*, simulating existing HD maps by removing pedestrian crossings and lane dividers, while retaining only the road boundaries.